OpenStudy (anonymous):

How do write cos(3x) in term of sin(x) ?

OpenStudy (anonymous):

By deMoivre formula I am able to find \[\cos 3x=4\cos^3 x - 3 cosx\] But I'm still not sure how to change it to sin(x) form.

OpenStudy (anonymous):

I don't understand why you people are taking on calculus when you don't even have the most basic of mathematical skills? You should have at least high school trigonometry to convert this, not deMoivre formulas or anything else. \[\cos x = \sin \left( \pi/2 + x \right)\]

OpenStudy (anonymous):

Hey. I found it out! Thx anyway. ;D Can you help me on my open question? This question is already closed.